OpenStudy (anonymous):

Which is an equation in point-slope form of the line that passes through the points (4,5) and (-3, -1)

OpenStudy (anonymous):

We need to find the slope first so \[\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

Can you find the slope using that formula please? ^_^

OpenStudy (anonymous):

\[11\frac{ ? }{ ? }-8\]

OpenStudy (anonymous):

oops sorry! 11 over -8

OpenStudy (anonymous):

Not quite I will set up the problem for you :) \[\huge~\rm~\frac{ -1-5 }{ -3-4 }=slope\]

OpenStudy (anonymous):

-6 over -7?

OpenStudy (anonymous):

yes :) Alright next step is to pick a point lets got with (4,5) Are you fimilar with this formula? \[\huge~\rm~y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

\[\huge~\rm~y-y_1=m(x-x_1)\] M=slope=6/7<--lets keep it positive : ) x1=4 y1=5 Can you sub in the values please? :)

OpenStudy (anonymous):

5-5=6/7(4-4)

OpenStudy (anonymous):

No quite we do not need to plug in the values for the varibles that have no number \[\huge~\rm~y-y_1=m(x-x_1)\] M=slope=6/7<--lets keep it positive : ) x1=4 y1=5 \[\huge~\rm~y-5=\frac{ 6 }{ 7 }(x-4)\]

OpenStudy (anonymous):

i suck at this stuff!

OpenStudy (anonymous):

Thats all it is buddy! :) You are doing just fine

OpenStudy (anonymous):

is that the answer? its not one of my choices

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

in here! sorry!

pooja195 (pooja195):

I can finish up :) looks like they left

OpenStudy (anonymous):

y+1=6/5(x+3) y+1=6/7(x+3) y-1=6/7(x-3) y+3=7/6(x+1)

pooja195 (pooja195):

oh ok the only problem is that they used the wrong point not hard to fix we always need to use the 2nd point so (-3,-1) and we get y+1=6/7(x+3) We get "+" because a negative and a negative touching each other make a postive :) understood?

OpenStudy (anonymous):

i do! thank you! could you help me with a few more?

OpenStudy (anonymous):

linear inequality graph Which is an equation of the given line in standard form? A. –11x + 6y = –13 B. –6x +11y = 13 C. –6x + 11y = –13 D. –6x + 7y = –17

pooja195 (pooja195):

wow i suck at standered for @Jaynator495

jaynator495 (jaynator495):

Cannot help because you are from k12, sorry about this... you can ask indirect questions but NO test questions...